OpenStudy (anonymous):

Erin recycles milk bottles, soda cans, and newspapers. If the following items are in the trash bin, what percent of the items can be recycled? 4 milk bottles, 1 banana peel, 6 newspapers, 4 soda cans, 5 apple cores 70% 100% 50% 14%

OpenStudy (anonymous):

@GracieBugg Hey can you help me out here? haha c:

OpenStudy (anonymous):

is it 50?

OpenStudy (graciebugg):

Give me a few seconds

OpenStudy (graciebugg):

Ok so I don't know it but Im tagging some smart people for you @SyedMohammed98 @undeadknight26 @bibby

OpenStudy (bibby):

the probability is items that can be recycled / total items how many items can be recycled?

OpenStudy (graciebugg):

Can I just give the answer?

OpenStudy (bibby):

I don't really care, other people do

OpenStudy (graciebugg):

Its 70% Because adding all of the items up equals 20 and 6 of them can't be recycled.6 is what percent of 20?Its 30%.So 100%-30%=70% so Its 70 hope that helped

OpenStudy (bibby):

4 milk bottles, 1 banana peel, 6 newspapers, 4 soda cans, 5 apple cores \(\dfrac{milk bottles, soda cans, and newspapers}{4+1+6+4+5}=\dfrac{14}{20}\)